OpenStudy (anonymous):

If sinA=5/6, where angle A is in Quadrant 1, find the exact value of sin (1/2A)

OpenStudy (notamathgenius):

1,172cm(2)

OpenStudy (noelgreco):

I assume that's sin (.5A). There is a half-angle formula for sin.

OpenStudy (anonymous):

so can you please help me solve the problem?

OpenStudy (ajprincess):

\[\cos^2A+\sin^2A=1\]\[\cos A=\sqrt{1-\sin^2A}\] From this \(\cos A\) can be found u can find \(\cos A\). Then use \[\cos A=1-2\sin^2(\frac{ 1 }{ 2 }A)\]\[\sin \frac{ 1 }{ 2 }A=\sqrt{\frac{ 1-\cos A }{ 2 }}\] from this u can find \[\sin \frac{ 1 }{2}A\] Does that help?

OpenStudy (anonymous):

but do u have an answer?

OpenStudy (notamathgenius):

Mine is at the very top

OpenStudy (anonymous):

did u use the half-angle formula because i'm getting a radical?

OpenStudy (ajprincess):

what r u getting?

OpenStudy (anonymous):